OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 16 <= 11 and x + 16 >= 3 11 >= x + 16 >= 3 -5 >= x >= -13 A number line with a closed circle on -13, a closed circle on -5, and shading in between. A number line with a closed circle on -13, shading to the left, and a closed circle on -5, shading to the right.

OpenStudy (anonymous):

x + 16 <= 11 x<=11-16 x<=-5 ================ x + 16 >= 3 x>=3-16 x>=-13

OpenStudy (anonymous):

Which of the Following is NOT equivalent?